OpenStudy (anonymous):

Multiply write your anwser in lowest terms 〖3z/8〗^7*3/z

OpenStudy (anonymous):

\[(\frac{3z}{8})^7\times \frac{3}{z}\]?

OpenStudy (anonymous):

do they really want you to find \[8^7\]?

OpenStudy (anonymous):

or just write in exponential notation?

OpenStudy (anonymous):

it wants the simplified product